what mass (in grams) of strontium phosphate is formed from the reaction of 29.0 mL of 0.0520 M SrCl2 solution with 12.0 mL of 0.110 M Na3PO4 Solution?
1
Expert's answer
2020-03-11T09:06:54-0400

What mass (in grams) of strontium phosphate is formed from the reaction of 29.0 mL of 0.0520 M SrCl2 solution with 12.0 mL of 0.110 M Na3PO4 Solution?


Solution:

The equation of the chemical reaction is following:

3SrCl2 + 2Na3PO4 = Sr3(PO4)2 + 6NaCl

We can calculate the mass of SrCl2 using given values:

m(SrCl2) = Cm(SrCl2) * M(SrCl2)*V = 0.0520 mol/L * (87.62+2*35.5) g/mol*0.029 L = 0.2392 (g)

The mass of Na3PO4 equals:

m(Na3PO4) = Cm(Na3PO4) * M(Na3PO4)*V = 0.110 mol/L * (3*23+1*31+4*16) g/mol*0.012 L = 0.2165 (g)

We can determine which of these reagents given in excess according to chemical equation:

0.2392 g X g

3SrCl2 + 2Na3PO4 = Sr3(PO4)2 + 6NaCl

3*158.62 2*164

Make a proportion:

0.2392/475.86 = X/328

where:

X = m(Na3PO4) = 0.2392*328/475.86 = 0.1649 (g)

Thus, Na3PO4 was given in excess and SrCl2 was given in lack.

So we can calculate the mass of product Sr3(PO4)2 using the mass of SrCl2:

0.2392 g X g

3SrCl2 + 2Na3PO4 = Sr3(PO4)2 + 6NaCl

3*158.62 452.86

Make a proportion:

0.2392/475.86 = X/452.86

where:

X = m(Sr3(PO4)2) = 0.2392*452.86/475.86 = 0.2276 (g)

Answer: m(Sr3(PO4)2) = 0.2276 g.